# Energy Window Augmented Plane Waves Approach to Density Functional Theory

### v6

@@ Abstract
In this work we present a new method for basis set generation for electronic structure calculations of crystalline solids. This procedure is aimed at applications to Density Functional Theory (DFT). In this construction, Energy Window Augmented Plane Waves (EWAPW), we take advantage of the fact that most DFT calculations use a convergence loop in order to obtain the self consistent eigenstates of the final (converged) Kohn Sham (KS) Hamiltonian. Here we propose that, for the basis used at each step of the self consistency iteration, we use the previous eigenstate basis, in the interstitial region, and augment it, inside each Muffin Tin (MT) sphere, with the solution to the spherically averaged KS Hamiltonian for the linearization energy of the energy window which contains the energy of that previous eigenstate. Indeed, to reduce the number of times the spherically averaged KS potential needs to be solved inside the MT spheres it is advantageous break up the spectrum into non-overlapping intervals, windows, and solve the spherically averaged KS Hamiltonian inside the MT region only once per window per angular momentum channel (at the linearization energy relevant to that window, usually near the middle of the window). For practical applications it is reasonable to have on the order of five to fifty windows. At each step of the iteration of the solution of the KS equations the EWAPW basis is that of near eigenstates of the KS Hamiltonian for that iteration. Overall the basis size is the comparable with the Augmented Plane Waves (APW) basis set but the number of radial wavefunctions is comparable or greater to Linearized Augmented Plane Waves + Local Orbitals + Higher Derivative Local Orbitals + High Energy Local Orbitals (LAPW+LO+HDLO+HELO).
@@ I Introduction
One solution to this glueing problem is the Linearized Augmented Plane Waves (LAPW) basis set [4, 2, 1]. Here plane waves are augmented within the MT sphere by a linear combination of the solutions to the spherically averaged KS Hamiltonian (see Eq. (1)) at the linearization energy $E_{L}^{n}$ (typically the middle of the valence band) $u_{l\alpha}^{n}\left(r,E_{L}^{n}\right)$ and its derivative with respect to energy $\dot{u}_{l\alpha}^{n}\left(r,E_{L}^{n}\right)=\frac{\partial}{\partial E}u_{l\alpha}^{n}\left(r,E_{L}^{n}\right)$. This is done for all angular momentum channels $lm$ (up to $l_{max}$) and all atoms $\alpha$. Whereby, by matching the MT wavefunctions to the planewave wavefunction, that is demanding the total function be continuous and continuously differentiable - at the MT sphere radius [4, 2, 1, 5, 6] - we obtain a efficient smooth basis set. The LAPW basis set may be written as:
@@
for $\nu=j$. One then truncates the basis set with number of basis functions $j\ll N_{\mathbf{G}}$, as one has a basis of near eigenvectors in the IR, and runs the KS loop to convergence. This greatly reduces the basis set size needed for the iterative calculations after the first iteration [18]. These methods can be further improved as in Quadratically Augmented Plane Waves QAPW [20, 21] where $\ddot{u}_{l\mu}^{n}\left(r,E_{L}^{n}\right)=\frac{\partial^{2}}{\partial E^{2}}u_{l\mu}^{n}\left(r,E_{L}^{n}\right)$ is also used inside the MT sphere and one more derivative (with respect to the radial co-ordinate) is matched to the plane waves in the IR which reduces linearization errors. However, the more derivatives are matched the more the radial wavefunction inside the MT sphere look like Bessel functions (indeed in the limit where an infinite number of derivatives is matched the radial wavefunction must be exactly Bessel) the higher cutoff, the harder the basis becomes [7, 18]. Furthermore if there are several relevant bands one can use Local Orbitals (LO) basis wave functions [8, 1, 9, 23, 18, 7]. Many LO like basis wavefunctions are of the form:
@@
In this work we will not be proposing any new type of LO (such as HDLO or HELO or conventional LO). As a matter of fact Energy Window Augmented Plane Wave (EWAPW) can be viewed as a competitor to the various forms of LO and lo, as we argue that LO and lo type wavefunctions will not significantly improve the accuracy of the basis set to obtain the eigenstates of the KS Hamiltonian (see Section III.6). Indeed, all of our basis, EWAPW, wavefunctions will be of the form $\frac{1}{\sqrt{\Omega}}\sum_{\mathbf{G}}\bar{o}_{\mathbf{k}\mathbf{G}}^{nj}\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}\right)$ in the IR, for appropriate $\bar{o}_{\mathbf{k}\mathbf{G}}^{nj}$ as in $\left(LA\right)^{2}PW$. However, unlike $\left(LA\right)^{2}PW$ methods, inside the MT region we will also be modifying the linearization energy of the $j$’th basis element based on the eigenvalue information from the previous, $n-1$’st, iteration of the solution of the KS equations. More precisely $\bar{o}_{\mathbf{k}\mathbf{G}}^{nj}$ will be chosen so that in the IR the basis wavefunctions have the same plane wave expansion as eigenstates $\nu$ of the previous iterations with $\nu=j$. In the MT we will be augmenting these wavefunction at the energies $E_{j}^{n}$ which are nearly equal to the eigenenergies of the previous iteration $\varepsilon_{\mathbf{k}\nu}^{n-1}$ (with $j=\nu$) except we will be using a widowing function $\mathcal{E}_{w}^{n}\left(E\right)$ (which is a piecewise constant approximation to the identity function: $Id\left(E\right)=E$, see Fig. (1)) and linearizing (augmenting) at $E_{j}^{n}=\mathcal{E}_{w}^{n}\left(\varepsilon_{\mathbf{k}\nu}^{n-1}\right)\cong\varepsilon_{\mathbf{k}\nu}^{n-1}$. As such one needs solve the spherically averaged KS Hamiltonian at a discrete (on the order of five to fifty) number of values $\mathcal{E}_{w}^{n}\left(E\right)$ takes on - per angular momentum channel, making it practical. Furthermore as we are able to have many linearization energies the method is likely more accurate then APW+lo or LAPW (comparable to LAPW+LO+HDLO+HELO), however we do not increase the basis size beyond APW levels. Detailed derivations are given in Section II below. We now note that the numerical efficiency of a DFT calculation is often determined by the diagonalization time and the time it takes to numerically set up the KS Hamiltonian and charge density [34]. Overall EWAPW has roughly the same diagonalization time as APW (same size of basis) but takes a comparable amount of time to set up the charge density and KS Hamiltonian as LAPW+LO+HDLO+HELO as it has a comparable number of radial basis functions. We note that in many situations with a large basis set the diagonalization time is the biggest [34].
@@ II.1 Wavefunctions
We now describe in detail the wavefunctions used in the EWAPW basis. EWAPW is a method to solve the question of which energy $E$ can be used to glue the wavefunction $u_{l\alpha}^{n}\left(E\right)$ in the MT to which combination of plane waves in the IR. This is done using information from the previous iteration of the solutions to the KS equations. We will not be using and LO/lo (or HDLO or HELO etc.) wavefunctions and the basis set size per $\mathbf{k}$ point will be the same as the number of reciprocal lattice vectors $N_{\mathbf{G}}$ (or less see Section III.4). We will also avoid solving for the radially averaged KS equation (see Eq. (1)) at every energy $\varepsilon_{\mathbf{k}\nu}^{n}$ (which is impractically hard) by using a windowing function $\mathcal{E}_{w}^{n}\left(E\right)$ that maps nearly via the identity into a small discrete set of values $\varepsilon_{\mathbf{k}\nu}^{n}\rightarrow E_{i,n}^{V}$ (see Fig. (1)) and only solving Eq. (1) at the energy values $E_{i,n}^{V}$.
We begin with a rigorous definition of a windowing function $\mathcal{E}_{w}^{n}\left(E\right)$, which is given by:
@@
Where $E_{1,n}^{U}=-\infty$, $E_{N,n}^{U}=+\infty$ and $E_{i-1,n}^{U}<E_{i,n}^{V}<E_{i,n}^{U}$ $\forall i$. Here $\Theta$ is the Heaviside function. Here $N-1$ , on the order of five to fifty, is the number of windows. The windowing function is a piecewise constant approximation to the identity function $Id\left(E\right)=E$. A typical $\mathcal{E}_{w}^{n}\left(E\right)$ is pictured in Fig. (1). We note that, as we shall see below, we will only need to solve Eq. (1) for the energies $E_{i,n}^{V}$. We further note that the windowing function divides the spectrum into non-overlapping intervals, windows, $E_{i-1,n}^{U}\leq E<E_{i,n}^{U}$ and assigns each window the value $E\rightarrow E_{i,n}^{V}$. Some methods to obtain $E_{i,n}^{V}$ and $E_{i,n}^{U}$ are described in Section (III.7). Here we simply note that, in general, it is advantageous to place many $E_{i,n}^{U}$ in the energy range of the valence and semicore bands (as such approximating the identity more accurately in those crucial regions) and a few for the rest of the spectrum (which is less crucial). We now introduce the wavefunctions: $\Phi_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)=$
@@
$\Psi_{\mathbf{k}\nu}^{n}=\sum_{\mathbf{G}}\bar{z}_{\mathbf{k}\mathbf{G}}^{\nu n}\Phi_{\mathbf{k}\mathbf{G}}^{n}\left(E^{n}\left(\nu,\mathbf{k},\mathbf{G}\right)\right)$ (14)
for some energies $E^{n}\left(\nu,\mathbf{k},\mathbf{G}\right)$. Then we that the EWAPW basis for the $n+1$’st iteration of the solution of the KS equations is given by:
@@
When $\mathcal{E}_{w}^{n+1}$ is nearly the identity function $\mathcal{E}_{w}^{n+1}\left(\varepsilon_{\mathbf{k}\nu}^{n}\right)\cong\varepsilon_{\mathbf{k}\nu}^{n}$ and the calculations are well converged $\varepsilon_{\mathbf{k}\nu}^{n+1}\cong\varepsilon_{\mathbf{k}\nu}^{n}$ this error is small, indeed there are multiple $E_{i,n}^{V}$ in the valence and semicore bands so the error $\left|\varepsilon_{\mathbf{k}\nu}^{n}-\mathcal{E}_{w}^{n+1}\left(\varepsilon_{\mathbf{k}\nu}^{n}\right)\right|$ is much smaller then bandwidth. We note that given that overall we have created a basis that is given by the previous eigenbasis in the interstitial and augmented it with the eigenstates of the spherically averaged KS Hamiltonian (with the linearization energy relevant of the energies of the previous basis, nearly $\varepsilon_{\mathbf{k}\nu}^{n}$ in the limit of many windows) with good convergence between iterations and good spherical symmetry in the MT spheres, in EWAPW, we are working with a basis of near eigenvectors. As such the KS Hamiltonian and overlap matrices will be nearly diagonal and sparse and as such quicker to diagonalize then the LAPW Hamiltonian and overlap matrices. We further note that unlike LAPW or APW+lo we have many radial wavefunctions $u_{l\alpha}^{n}\left(r,E_{i,n}^{V}\right)$ per angular momentum channel, it is a comparable or greater amount of radial wavefunctions to LAPW+LO+HDLO+HELO. Furthermore we note that we only have $N_{\mathbf{G}}$ basis wavefunctions per $\mathbf{k}$ point. Therefore it is reasonable to surmise that we have a basis with a comparable diagonalization speed to APW and comparable accuracy to LAPW+LO+HDLO+HELO. We note that setting up the KS Hamiltonian, overlap and charge density will take comparable time to LAPW+LO+HDLO+HELO methods.
@@ III.2 Partial Diagonalization
In many situations it is advantageous for computational speed to only look for the lowest few bands and not diagonalize the entire KS Hamiltonian matrix [1]. Indeed the higher bands are unoccupied so do not contribute to the electron density and total energy. In which case since we are interested in only the lowest bands, say only $\mathcal{M_{D}}$ of them have the energies $\varepsilon_{\mathbf{k}\nu}^{n},\bar{z}_{\mathbf{k}\mathbf{G}}^{\nu n}$ computed with $\nu\leq\mathcal{M_{D}}$. Then for the first $\mathcal{M_{D}}$ bands we proceed as in Section II.1. The rest of the basis may be obtained as in Eq. (21) with $0\leftrightarrow n$ and the $\mathbf{G}$’s being chosen such that the smallest $\mathcal{M}$ energies in Eq. (21) are skipped. That is for the higher $N_{\mathbf{G}}-\mathcal{M_{D}}$ bands we write:
@@ III.4 Partial basis (compatibility with $BEE-\left(LA\right)^{2}PW$ ideas [18])
We note that the number of basis elements $\Psi_{\mathbf{k}j}^{n+1}$ may be significantly smaller then the value $N_{\mathbf{G}}$ of reciprocal wavevectors. Indeed once the system has been initialized it is a basis of near eigenvectors it is permissible to significantly reduce basis size. Indeed, as we have a basis of near eigenvectors the Hamiltonian and overlap matrices are nearly diagonal and sparse. As such it is permissible to truncate the basis at much smaller basis numbers then $N_{\mathbf{G}}$. We note that this means that in the interstitial we limit ourselves to a smaller basis set but one that is adapted to an earlier IR for an earlier KS Hamiltonian. To overcome this last minor difficulty [18] we can periodically (every several iterations of the solutions of the KS equations) expand the basis as in Section III.3 to increase the basis inside the interstitial while still saving on computer time overall. We will call this method BEE-EWAPW [18].
III.5 EWAPW vs. Energy Window Linearized Augmented Plane Waves (EWLAPW), multi-radius [26] and Continuous Interpolation Energy Window Augmented Plane Waves (CEWAPW).
The EWLAPW basis is described in Appendix A while multi radius options are described in Appendix B (see also ref. [26]) and CEWAPW is described in Appendix C. While these methods are slightly advantageous over EWAPW - in that they are slightly more accurate; the added complexity does not seem to be worth the minor improvement. To see this, we notice that the KS Hamiltonian, $-\nabla^{2}+V_{KS}^{n+1}\left(\mathbf{r}\right)$, has no singularities on the MT sphere radius. As such the eigenwavefunctions of the KS Hamiltonian should have no singularities near the surface of the MT sphere. Since the basis wavefunctions proposed in EWAPW are near eigenstates they should have no large singularities near the surface of the MT spheres and should be nearly continuously differentiable even though the constituent wavefunctions given in Eq. (11) have derivative discontinuities. As such adding $\dot{u}_{l\alpha}^{n}\left(r,\mathcal{E}_{w}^{n}\left(\varepsilon_{\mathbf{k}\nu}^{n}\right)\right)$ or varying the radius $S_{\alpha}$ to make the constituent wavefunctions continuously differentiable would not make much of a difference. From another perspective it is true that the scaling of the linearization error improves, we have that the linearization error changes:
@@
however it is already very small as the basis is nearly complete see Eq. (17) and the discussion below when the bands are well split into multiple windows so the improvement of multi radius, CEWAPW and EWLAPW options is minor. We note that continuous interpolation, multi radius and linearization options are combinable with each other.
III.6 LO and lo and related extensions
We note that LO and lo extensions as well as HDLO and HELO are not essential. The basis set already has a large number of wavefunctions at various energies of the spherically averaged KS Hamiltonian inside the MT spheres so an additional wavefunction with energy close to those used in the construction of the EWAPW basis would make limited difference - there are already enough radial wavefunction in the EWAPW basis. This increase of basis set number is not essential for accuracy but is quite detrimental to computational speed (see Section V).
III.7 Generating the energy windows

### v7

@@ Abstract
In this work we present a new method for basis set generation for electronic structure calculations of crystalline solids. This procedure is aimed at applications to Density Functional Theory (DFT). In this construction, Energy Window Augmented Plane Waves (EWAPW), we take advantage of the fact that most DFT calculations use a convergence loop in order to obtain the self consistent eigenstates of the final (converged) Kohn Sham (KS) Hamiltonian. Here we propose that, for the basis used at each step of the self consistency iteration, we use the previous eigenstate basis, in the interstitial region, and augment it, inside each Muffin Tin (MT) sphere, with the solution to the spherically averaged KS Hamiltonian for the linearization energy of the energy window which contains the energy of that previous eigenstate. Indeed, to reduce the number of times the spherically averaged KS potential needs to be solved inside the MT spheres it is advantageous break up the spectrum into non-overlapping intervals, windows, and solve the spherically averaged KS Hamiltonian inside the MT region only once per window per angular momentum channel (at the linearization energy relevant to that window, usually near the middle of the window). For practical applications it is reasonable to have on the order of five to fifty windows. At each step of the iteration of the solution of the KS equations the EWAPW basis is that of near eigenstates of the KS Hamiltonian for that iteration. Overall the basis size is the comparable with the Augmented Plane Waves (APW) basis set but the number of radial wavefunctions is comparable or greater to Linearized Augmented Plane Waves + Local Orbitals + Higher Derivative Local Orbitals + High Energy Local Orbitals (LAPW+LO+HDLO+HELO) basis set.
@@ I Introduction
Another solution to this glueing problem is the Linearized Augmented Plane Waves (LAPW) basis set [4, 2, 1]. Here plane waves are augmented within the MT sphere by a linear combination of the solutions to the spherically averaged KS Hamiltonian (see Eq. (1)) at the linearization energy $E_{L}^{n}$ (typically the middle of the valence band) $u_{l\alpha}^{n}\left(r,E_{L}^{n}\right)$ and its derivative with respect to energy $\dot{u}_{l\alpha}^{n}\left(r,E_{L}^{n}\right)=\frac{\partial}{\partial E}u_{l\alpha}^{n}\left(r,E_{L}^{n}\right)$. This is done for all angular momentum channels $lm$ (up to $l_{max}^{L}$) and all atoms $\alpha$. Whereby, by matching the MT wavefunctions to the planewave wavefunction, that is demanding the total function be continuous and continuously differentiable - at the MT sphere radius [4, 2, 1, 5, 6] - we obtain a efficient continuously differentiable basis set. The LAPW basis set may be written as:
@@
for $\nu=j$. One then truncates the basis set with number of basis functions $j\ll N_{\mathbf{G}}$ (where $N_{\mathbf{G}}$ is the number of reciprocal lattice points), as one has a basis of near eigenvectors in the IR, and runs the KS loop to convergence. This greatly reduces the basis set size needed for the iterative calculations after the first iteration [19]. These methods can be further improved as in Quadratically Augmented Plane Waves QAPW [21, 22] where $\ddot{u}_{l\mu}^{n}\left(r,E_{L}^{n}\right)=\frac{\partial^{2}}{\partial E^{2}}u_{l\mu}^{n}\left(r,E_{L}^{n}\right)$ is also used inside the MT sphere and one more derivative (with respect to the radial co-ordinate) is matched to the plane waves in the IR which reduces linearization errors. However, the more derivatives are matched the more the radial wavefunction inside the MT sphere look like Bessel functions (indeed in the limit where an infinite number of derivatives is matched the radial wavefunction must be exactly Bessel) the higher cutoff, the harder the basis becomes [7, 19]. Furthermore if there are several relevant bands one can use Local Orbitals (LO) basis wave functions [8, 1, 9, 24, 19, 7]. Many LO like basis wavefunctions are of the form:
@@
In this work we will not be proposing any new type of LO (such as HDLO or HELO or conventional LO). As a matter of fact Energy Window Augmented Plane Wave (EWAPW) can be viewed as a competitor to the various forms of LO and lo, as we argue that LO and lo type wavefunctions will not significantly improve the accuracy of the basis set to obtain the eigenstates of the KS Hamiltonian (see Section III.7). Indeed, all of our basis, EWAPW, wavefunctions will be of the form $\frac{1}{\sqrt{\Omega}}\sum_{\mathbf{G}}\bar{o}_{\mathbf{k}\mathbf{G}}^{nj}\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}\right)$ in the IR, for appropriate $\bar{o}_{\mathbf{k}\mathbf{G}}^{nj}$ as in $\left(LA\right)^{2}PW$. However, unlike $\left(LA\right)^{2}PW$ methods, inside the MT region we will also be modifying the linearization energy of the $j$’th basis element based on the eigenvalue information from the previous, $n-1$’st, iteration of the solution of the KS equations. More precisely $\bar{o}_{\mathbf{k}\mathbf{G}}^{nj}$ will be chosen so that in the IR the basis wavefunctions have the same plane wave expansion as eigenstates $\nu$ of the previous iterations with $\nu=j$. In the MT we will be augmenting these wavefunction at the energies $E_{j}^{n}$ which are nearly equal to the eigenenergies of the previous iteration $\varepsilon_{\mathbf{k}\nu}^{n-1}$ (with $j=\nu$) except we will be using a widowing function $\mathcal{E}_{w}^{n}\left(E\right)$ (which is a piecewise constant approximation to the identity function: $Id\left(E\right)=E$, see Fig. (1)) and linearizing (augmenting) at $E_{j}^{n}=\mathcal{E}_{w}^{n}\left(\varepsilon_{\mathbf{k}\nu}^{n-1}\right)\cong\varepsilon_{\mathbf{k}\nu}^{n-1}$. As such one needs solve the spherically averaged KS Hamiltonian at a discrete (on the order of five to fifty) number of values $\mathcal{E}_{w}^{n}\left(E\right)$ takes on - per angular momentum channel, making it practical. Furthermore as we are able to have many linearization energies the method is likely more accurate then APW+lo or LAPW (comparable to LAPW+LO+HDLO+HELO), however we do not increase the basis size beyond APW levels. Detailed derivations are given in Section II below. We now note that the numerical efficiency of a DFT calculation is often determined by the diagonalization time and the time it takes to numerically set up the KS Hamiltonian and charge density [29]. Overall EWAPW has roughly the same diagonalization time as APW (same size of basis) but takes a comparable amount of time to set up the charge density and KS Hamiltonian as LAPW+LO+HDLO+HELO as it has a comparable number of radial basis functions. We note that in many situations with a large basis set the diagonalization time is the biggest [29].
@@ II.1 Wavefunctions
We now describe in detail the wavefunctions used in the EWAPW basis. EWAPW is a method to solve the question of which energy $E$ can be used to glue the wavefunction $u_{l\alpha}^{n}\left(E\right)$ in the MT to which combination of plane waves in the IR. This is done using information from the previous iteration of the solutions to the KS equations. We will not be using and LO/lo (or HDLO or HELO etc.) wavefunctions and the basis set size, per $\mathbf{k}$ point, will be the same as the number of reciprocal lattice vectors $N_{\mathbf{G}}$ (or less see Section III.4). We will also avoid solving for the radially averaged KS equation (see Eq. (1)) at every energy $\varepsilon_{\mathbf{k}\nu}^{n}$ (which is impractically hard) by using a windowing function $\mathcal{E}_{w}^{n}\left(E\right)$ that maps nearly via the identity into a small discrete set of values $\varepsilon_{\mathbf{k}\nu}^{n}\rightarrow E_{i,n}^{V}$ (see Fig. (1)) and only solving Eq. (1) at the energy values $E_{i,n}^{V}$.
We begin with a definition of a windowing function $\mathcal{E}_{w}^{n}\left(E\right)$, which is given by:
@@
Where $E_{1,n}^{U}=-\infty$, $E_{N,n}^{U}=+\infty$ and $E_{i-1,n}^{U}<E_{i,n}^{V}<E_{i,n}^{U}$ $\forall i$. Here $\Theta$ is the Heaviside function. Here $N-1$ , on the order of five to fifty, is the number of windows. The windowing function is a piecewise constant approximation to the identity function $Id\left(E\right)=E$. A typical $\mathcal{E}_{w}^{n}\left(E\right)$ is pictured in Fig. (1). We note that, as we shall see below, we will only need to solve Eq. (1) for the energies $E_{i,n}^{V}$. We further note that the windowing function divides the spectrum into non-overlapping intervals, windows, $E_{i-1,n}^{U}\leq E<E_{i,n}^{U}$ and assigns each window the value $E\rightarrow E_{i,n}^{V}$. Some methods to obtain $E_{i,n}^{V}$ and $E_{i,n}^{U}$ are described in Section (III.8). Here we simply note that, in general, it is advantageous to place many $E_{i,n}^{U}$ in the energy range of the valence and semicore bands (as such approximating the identity more accurately in those crucial regions) and a few for the rest of the spectrum (which is less crucial). We now introduce the wavefunctions: $\Phi_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)=$
@@
$\Psi_{\mathbf{k}\nu}^{n}\left(\mathbf{r}\right)=\left\{\begin{array}[]{cc}\sum_{\mathbf{G}}\bar{z}_{\mathbf{k}\mathbf{G}}^{\nu n}\frac{1}{\sqrt{\Omega}}\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}\right)&IR\\ \mathrm{Arbitrary}&MT\end{array}\right.$ (14)
that is we focus on the expansion of the wavefunction in the IR. Then we that the EWAPW basis for the $n+1$’st iteration of the solution of the KS equations is given by:
@@
When $\mathcal{E}_{w}^{n+1}$ is nearly the identity function $\mathcal{E}_{w}^{n+1}\left(\varepsilon_{\mathbf{k}\nu}^{n}\right)\cong\varepsilon_{\mathbf{k}\nu}^{n}$ and the calculations are well converged $\varepsilon_{\mathbf{k}\nu}^{n+1}\cong\varepsilon_{\mathbf{k}\nu}^{n}$ this error is small, indeed there are multiple $E_{i,n}^{V}$ in the valence and semicore bands so the error $\left|\varepsilon_{\mathbf{k}\nu}^{n}-\mathcal{E}_{w}^{n+1}\left(\varepsilon_{\mathbf{k}\nu}^{n}\right)\right|$ is much smaller then bandwidth. We note that given that overall we have created a basis that is given by the previous eigenbasis in the interstitial and augmented it with the eigenstates of the spherically averaged KS Hamiltonian (with the linearization energy relevant of the energies of the previous basis, nearly $\varepsilon_{\mathbf{k}\nu}^{n}$ in the limit of many windows) with good convergence between iterations and good spherical symmetry in the MT spheres, in EWAPW, we are working with a basis of near eigenvectors. As such the KS Hamiltonian and overlap matrices will be nearly diagonal and sparse and as such quicker to diagonalize then the LAPW Hamiltonian and overlap matrices. We further note that unlike LAPW or APW+lo we have many radial wavefunctions $u_{l\alpha}^{n}\left(r,E_{i,n}^{V}\right)$ per angular momentum channel, it is a comparable or greater amount of radial wavefunctions to LAPW+LO+HDLO+HELO. Furthermore we note that we only have $N_{\mathbf{G}}$ basis wavefunctions per $\mathbf{k}$ point. Therefore it is reasonable to surmise that we have a basis with a comparable diagonalization speed to APW and comparable accuracy to LAPW+LO+HDLO+HELO. We note that setting up the KS Hamiltonian, overlap and charge density will take comparable time to setup as in LAPW+LO+HDLO+HELO methods.
@@ III.2 Partial Diagonalization
In many situations it is advantageous for computational speed to only look for the lowest few bands and not diagonalize the entire KS Hamiltonian matrix [1]. Indeed the higher bands are unoccupied so do not contribute to the electron density and total energy. In which case since we are interested in only the lowest bands, say only $\mathcal{M_{D}}$ of them, we have that the energies $\varepsilon_{\mathbf{k}\nu}^{n},\bar{z}_{\mathbf{k}\mathbf{G}}^{\nu n}$ are computed only with $\nu\leq\mathcal{M_{D}}$. Then for the first $\mathcal{M_{D}}$ bands we proceed as in Section II.1. The rest of the basis may be obtained as in Eq. (21) with $0\leftrightarrow n$ and the $\mathbf{G}$’s being chosen such that the smallest $\mathcal{M}$ energies in Eq. (21) are skipped. That is for the higher $N_{\mathbf{G}}-\mathcal{M_{D}}$ bands we write:
@@ III.4 Partial basis (compatibility with $BEE-\left(LA\right)^{2}PW$ ideas [19])
Here we note that the number of basis elements $\Psi_{\mathbf{k}j}^{n+1}$ may be significantly smaller then the value $N_{\mathbf{G}}$ of reciprocal wavevectors. Indeed once the system has been initialized it is a basis of near eigenvectors it is permissible to significantly reduce basis size. Indeed, as we have a basis of near eigenvectors the Hamiltonian and overlap matrices are nearly diagonal and sparse. As such it is permissible to truncate the basis at much smaller basis numbers then $N_{\mathbf{G}}$. We note that this means that in the interstitial we limit ourselves to a smaller basis set but one that is adapted to an earlier IR for an earlier KS Hamiltonian. In principle, to overcome this last minor difficulty [19] we can periodically (every several iterations of the solutions of the KS equations) expand the basis as in Section III.3 to increase the basis inside the interstitial while still saving on computer time overall. We will call this method BEE-EWAPW [19].
III.5 EWAPW vs. Energy Window Linearized Augmented Plane Waves (EWLAPW), multi-radius [30] options.
The EWLAPW basis is described in Appendix A while multi radius options are described in Appendix B (see also ref. [30]). While these methods are slightly advantageous over EWAPW - in that they are slightly more accurate; the added complexity does not seem to be worth the minor improvement. To see this, we notice that the KS Hamiltonian, $-\nabla^{2}+V_{KS}^{n+1}\left(\mathbf{r}\right)$, has no singularities on the MT sphere radius. As such the eigenwavefunctions of the KS Hamiltonian should have no singularities near the surface of the MT sphere. Since the basis wavefunctions proposed in EWAPW are near eigenstates they should have no large singularities near the surface of the MT spheres and should be nearly continuously differentiable even though the constituent wavefunctions given in Eq. (11) have derivative discontinuities. As such adding $\dot{u}_{l\alpha}^{n}\left(r,\mathcal{E}_{w}^{n}\left(\varepsilon_{\mathbf{k}\nu}^{n}\right)\right)$ or varying the radius $S_{\alpha}$ to make the constituent wavefunctions continuously differentiable would not make much of a difference. From another perspective it is true that the scaling of the linearization error improves, we have that the linearization error changes:
@@
however it is already very small as the basis is nearly complete see Eq. (17) and the discussion below when the bands are well split into multiple windows so the improvement of multi radius and EWLAPW options is minor.
III.6 Continuous Interpolation Energy Window Augmented Plane Waves (CEWAPW)
We note that CEWAPW is described in Appendix C. We note that CEWAPW extensions are not essential in the limit of a large number of windows but become more important for a moderate number of windows where they can enhance accuracy at a limited computational cost. We note that continuous interpolation, multi radius and linearization options are combinable with each other.
III.7 LO and lo and related extensions
We note that LO and lo extensions as well as HDLO and HELO are not essential. The basis set already has a large number of wavefunctions at various energies of the spherically averaged KS Hamiltonian inside the MT spheres so an additional wavefunction with energy close to those used in the construction of the EWAPW basis would make limited difference - there are already enough radial wavefunction in the EWAPW basis. In principle it is possible to remove some windows around the energies of the LO wavefunctions and use a larger basis involving LO however that seems disadvantageous. For example, one could remove all the semicore windows; focus on putting many windows in the valence band and add some LO, however there are limited advantages to this. Indeed, this increase of basis set number is not essential for accuracy but is quite detrimental to computational speed (see Section V).
III.8 Generating the energy windows